taylor puts $1000 in a savings account that pays 4% interest, compounded annually. how much money will be in the account 3 years later if she makes no more deposits.
Answer by JulietG(1812)

Compounding at the end of the first year -
1000 * 1.04 = 1040
Compounding at the end of the second year -
1040 * 1.04 = 1124.86
Compounding at the end of the third year -
1124.86 * 1.04 = 1169.85
This is how credit cards make their money, only they are compounding DAILY or continuously, so you're paying interest on your interest from the very beginning.